# pnasresearcharticle

### v1

@@ Discussion
Sangelose hydrogels were prepared by dissolving Sangelose 90L (Daido Chemical Co., Osaka, Japan) powder in phosphate-buffered saline (PBS) to form a 6% w/w stock, which was then diluted to 3% w/w. Tween 80 was added during dilution as needed. Detailed formulation procedures are provided in the SI.
@@
The diffusion coefficients of probe cargoes in the hydrogels were determined using fluorescence recovery after photobleaching (FRAP), procedure is detailed in the SI. Fluorescein isothiocyanate-tagged bovine serum albumin (BSA, Sigma, $M_{\rm w}\simeq 68\leavevmode\nobreak\ {\rm kDa}$) was used for PVA-Borax, and Alexa 647-tagged human immunoglobulin (hIgG, Sigma, $M_{\rm w}\simeq 150\leavevmode\nobreak\ {\rm kDa}$) for Sangelose.
Briefly, fluorescent cargo was bleached with a high-intensity laser, and the fluorescence recovery time constant, $t_{1/2}$, was used to calculate the diffusion coefficient $\mathbb{D}=\gamma_{\mathbb{D}}(\rho^{2}/t_{1/2})$, where $\gamma_{\mathbb{D}}$ is an empirically determined constant (37).
@@
Release assays were conducted by injecting $100\leavevmode\nobreak\ \mu{\rm L}$ of PVA-Borax and Sangelose hydrogels loaded with FITC-BSA and Alexa-hIgG respectively into a capillary tube filled with $400\leavevmode\nobreak\ \mu{\rm L}$ of the appropriate buffer. At regular intervals, the supernatant was removed and analyzed to measure the amount of cargo present. Detailed procedures are described in the SI.
@@
Statistical Analyses
The statistical analyses to determine the significance levels across datasets for the in vitro and in vivo data were performed using a Kolmogorov-Smirnov test in MATLAB with a 5% threshold.
[…]
Supplementary Information and Datasets
All experimental data shown in the figures are provided as .xls files in the Supplementary Datasets section.
[…]
@@
[…]
[40]

### v2

@@ Methods
The diffusion coefficients of probe cargoes in the hydrogels were determined using fluorescence recovery after photobleaching (FRAP), procedure is detailed in the SI. Fluorescein isothiocyanate-tagged bovine serum albumin (BSA, Sigma, $M_{\rm w}\simeq 68~{}{\rm kDa}$) was used for PVA-Borax, and Alexa 647-tagged human immunoglobulin (hIgG, Sigma, $M_{\rm w}\simeq 150~{}{\rm kDa}$) for Sangelose. Briefly, fluorescent cargo was bleached with a high-intensity laser, and the fluorescence recovery time constant, $t_{1/2}$, was used to calculate the diffusion coefficient $\mathbb{D}=\gamma_{\mathbb{D}}(\rho^{2}/t_{1/2})$, where $\gamma_{\mathbb{D}}$ is an empirically determined constant [4].
@@
Release assays were conducted by injecting $100~{}\mu{\rm L}$ of PVA-Borax and Sangelose hydrogels loaded with FITC-BSA and Alexa-hIgG respectively into a capillary tube filled with $400~{}\mu{\rm L}$ of the appropriate buffer. At regular intervals, the supernatant was removed and analyzed to measure the amount of cargo present. Detailed procedures are described in the SI.
@@
Resource Availability
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Eric A. Appel (eappel@stanford.edu).
Materials availability
This study did not generate any new materials.
Data and code availability
  • All experimental data shown in the figures are provided as .xls files in the Supplementary Datasets section. Detailed description of the aforementioned experimental techniques are provided in the Supplementary Information PDF.
  • All original code used to analyze the rheological viscoelastic spectrum data, diffusivity measurements, and pharmacokinetics modeling are available from the lead author upon reasonable request.
  • Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
Acknowledgments
@@
Author Contributions
S.S., E.A.A. conceived the idea. S.S., C.D., C.K.J., Y.E.S., A.N.P. performed experiments. S.S., C.K.J. analyzed the data. S.S., C.D., C.K.J., E.A.A. wrote the manuscript.
Declaration of Interests